# Exact Model Reduction for Continuous-Time Open Quantum Dynamics

### v2

@@ Abstract
I Introduction
I.1 Context and motivation
@@
Motivated by the above applications, a host of strategies for reducing the demands of direct simulation of quantum dynamics and obtaining tractable “effective” models have been developed. As in the classical case, a main guiding principle is the fact that, even if access to a complete description would be granted, this would neither be desirable nor necessary, as only certain information is ultimately “relevant” on physical grounds. Discarding “irrelevant” information can be achieved by invoking different forms of coarse-graining, aggregating, or averaging over microscopic degrees of freedom, typically in combination with perturbative arguments – paradigmatic example being renormalization-group approaches, mean-field and hydrodynamic descriptions, or cluster-expansion techniques [17, 18, 19]. For open quantum systems specifically, weak-coupling and timescale-separation assumptions are used in standard Born-Markov derivations of master equations [20, 2, 21], along with more sophisticated approaches based, for instance, on different forms of clustering or coarse graining in space [22, 23] or time [24, 25]. Further to that, acknowledging that the dynamics of realistic many-body quantum systems (subject to specified locality, energy, and time constraints) may a priori be confined to an exponentially small fraction of the available state space [26] in turn legitimates efficient parametrizations of “relevant” quantum states in terms of tensor networks [27, 19]. The use of tensor-network methods, in combination with chain-mapping techniques [28, 29] or process-tensor representations [30, 31], lies at the core of current state-of-the-art numerical methods available for open-system simulation in the presence of highly structured, possibly strongly-coupled environments. As the conceptual and practical significance of open quantum dynamics continues to expand across quantum science,efforts for obtaining improved simulation methods are multiplying, with a special focus on the simplest yet paradigmatic case of dynamics described by time-independent Lindblad master equations [20]. Recent representative contributions include more accurate and efficient classical simulation algorithms based on high-order adiabatic elimination [32, 33] or low-rank structure [34], as well as quantum simulation algorithms that exploit unitary embeddings [35] or classical randomness [36].
@@
From a system-theoretical standpoint, the different approaches mentioned above can all be seen as instances of approximate model (-order) reduction (MR), a general framework aimed at reducing the complexity of a mathematical model while retaining some of the essential features of the original one [37]. A familiar setting in which MR takes place, in fact in its strongest exact form, arises when the dynamics exhibits a manifest symmetry. If the dynamics are described by a Hamiltonian, a decomposition of the Hilbert space according to irreducible representations of the symmetry group naturally furnishes invariant subspaces [38]; hence, if the initial state resides entirely in one such subspace, an exact lower-dimensional Hamiltonian model may be found by restricting to this invariant subspace via projection. This holds true both for “conventional” symmetries, for which the number of dynamically disconnected invariant, “Krylov subspaces” [39, 40] grows at most polynomially with system size, and for non-manifest “unconventional” symmetries, as occurring in systems exhibiting so-called Hilbert space fragmentation [41, 42, 43], whereby exponentially many disconnected Krylov subspaces exist. In this context, an important step was taken by Kumar and Sarovar [44], where conditions for certifying the existence of a non-trivial proper invariant subspace were identified and, building on that, exact MR methods were developed for quantum many-body Hamiltonian models. A subspace of the Hilbert space, however, is not the natural candidate for reduction when the initial condition is not a pure state or the dynamics is dissipative. As such, the approach does not lend itself to the application to open quantum dynamics.
@@
Our framework may describe a many-body open quantum system $S$ evolving on a (finite-dimensional) Hilbert space $\mathcal{H}$ and subject to Markovian dissipation, or it can be applied to reduce the joint (Hamiltonian or Lindblad) dynamics of a many-body bipartite system evolving on $\mathcal{H}\equiv\mathcal{H}_{S}\otimes\mathcal{H}_{E}$, with $E$ being a quantum environment – in which case the reduced dynamics of $S$ alone is generically non-Markovian. While being suitable for describing arbitrary quantum states (including non-factorized system-environment initial conditions), our approach explicitly leverages the fact that, in typical applications, the dynamical behavior of the system is not only needed solely for certain initial conditions, but also for only certain observable properties of interest (e.g., expectations of few-body or collective observables, diagonal elements of the density operator, and so on). This results in the possibility of carrying out dimensional MR based on both initial conditions and output observables, which has not been exploited before to the best of our knowledge. By defining the dynamical models of interest on operator (*-)algebras [47], and insisting that projections are also taken onto operator subalgebras, we can guarantee that the reduced models we obtain obey the structural constraints of physically admissible (completely positive and trace-preserving, CPTP) quantum dynamics – which we prove are still in Lindblad form. An iterative algorithm is provided to construct a reduced, lower-dimensional Lindbladian model, when one exists. We further show that, even in situations where symmetries are present in the original Lindblad dynamics, our general framework may allow for more effective reduction than achievable by known symmetry-based approaches, through the identification of a new type of observable-dependent symmetries.
@@
Related work. From a technical standpoint, we borrow some key tools and results from our previous work on MR for classical hidden Markov models [48] and quantum discrete-time dynamics [49]. While our guiding philosophy here is similar, extension to continuous-time dynamics nonetheless entails, as we will see, several non-trivial extensions. In particular, it is by no means obvious – and it is in fact remarkable – that our MR procedures yield a valid Lindblad generator. Further to that, extension to continuous-time dynamics substantially broaden the class of physical settings and phenomena to which our methods may be usefully applied to, given the foundational importance of Lindblad dynamics.
@@
Concerning early contributions to the use of MR techniques for quantum dynamics, projection-based methods were employed for nonlinear conditional (stochastic) quantum master equations in the context of cavity QED [50], and later used to exactly derive a special (Maxwell-Bloch-type) class of Lindblad dynamics [51]. In [52], MR by balanced truncation was exploited to derive reduced (exact or approximate) models for linear quantum stochastic dynamics, with emphasis on providing conditions under which physical realizability constraints are preserved. Closer to our work, both in spirit and in terms of the algebraic formalism employed, are the work in [44] we mentioned above and the work in [53], where a notion of state-space coarse-graining is formulated for quantum systems. As we already noted, while the approach in [44] also achieves exact quantum MR, a quantum model is obtained “for free” because only Hamitonian and pure states are considered. On the flip side, while the approach of [53] is not a priori restricted to Hamiltonian dynamics, the coarse-grained reduced models are not always exact nor provably CPTP or in Lindblad form. More recently, building on the concept of “lumpability” of Markov processes [54], exact quantum MR techniques based on constrained bisimulation have been developed for boosting the simulation of quantum circuits [55]. Aside from the fact that this approach targets discrete-time dynamics (rather than continuous-time as we consider here), it is again restricted to unitary evolution and pure states in current form.
@@
I.2 Structure and summary of main results
@@
The dynamical models we focus on are formally introduced in Sec. II, in terms of a Lindblad master equation (in the Schrödinger or dual, Heisenberg picture) paired with a linear output equation – which we term a quantum-dynamical semigroup with output. After exemplifying some basic scenarios of interest this encompasses, we pose two dimensional MR problems, distinguished by the type of reduced model one is seeking: a reduced CPTP quantum dynamics or, respectively, a general linear dynamical system – in such a way that the specified outputs are exactly reproduced for all times. The essential idea we use for obtaining the desired reduction is also anticipated here: namely, the reduction of the dynamics to an appropriate Krylov operator subspace.
@@
The next two sections of the paper, Sec. III and Sec. IV, are devoted to introduce the mathematical apparatus needed for tackling and solving the MR problems we posed. In particular, the requisite Krylov operator subspaces are built by leveraging the available knowledge of the initial conditions of interest for the system (reachable subspaces, introduced in Sec. III.1) and the target output observables for the dynamics (observable subspaces, constructed in Sec. IV.1). In order to extend such subspaces to operator subalgebras, amenable to support quantum dynamics, some existing structural results on operator *-algebras and their associated quantum conditional expectations and “distortion maps” are needed; these are recalled in Sec. III.2. Our key results in these sections may be summarized as follows:
  • Linear reductions. The solution to the linear MR problem is derived in Theorem 1 based on the reachable subspace, and in Theorem 5 based on the observable subspace. Both solutions are provably optimal in the sense of having minimal dimension, and can be combined to find a minimal reduction by considering the effective subspace introduced in Sec. IV.3.
@@
  • Quantum reductions. The solution to the quantum MR problem is established in Theorem 4, where we prove that a semigroup generator with output retains its CPTP properties when appropriately restricted to an operator subalgebra, and propose a constructive algorithm to obtain the desired reduction starting from the specified set of initial conditions. Notably, the initial states can be restricted to an arbitrary linear set, without the need to assume initial factorization in the case a composite system-environment setting, as in standard derivations of master equations [2]. In Sec. IV.2, we derive an equivalent result for reductions based on the observable set of interest, and the two procedures are combined in an iterative algorithm described in Sec. IV.3.
@@
Section V bridges between the general theory sections and the illustrative applications that will follow, by focusing on the interplay of our general MR procedures with symmetries which, for Lindblad dynamics, are usually categorized as “strong” or “weak” – depending on whether, loosely speaking, invariance occurs at the level of individual (Hamiltonian and Lindblad) operators or only at the superoperator level [60, 61]. We show how viable subspaces and algebras for MR naturally emerge from considering symmetries, and characterize the sense in which a strong symmetry leads to a computational advantage over a weak one, from a dimensional-reduction standpoint. Nonetheless, focusing on observable-based MR, we prove in Theorem 7 that our methods can improve reductions based on both strong and weak symmetries. In fact, finding the relevant output algebra by the procedure described above corresponds to finding “generalized symmetries” that emerge only when considering the observables of interest: this new class of observable-dependent symmetries is defined in Sec. V.4.
@@
In Sec. VI, we examine concrete examples motivated by condensed-matter physics and quantum information science applications, with a twofold goal: on the one hand, to benchmark our general approach, by showing that it consistently recovers known exact results when available, or it yields the correct numerical results of the full model, as expected; on the other hand, to showcase additional flexibility and resource reduction, beyond what achievable with existing methods to the best of our knowledge. Specifically, in Sec. VI.1 we study variations of the well-known central-spin model, beginning with the case where the dynamics is Hamiltonian (no dissipation) and enjoys permutation symmetry on the bath spins. Here, we observe how the introduction of a bath Hamiltonian $H_{B}$ influences the achievable MR, by showing how, even when $H_{B}$ breaks permutation symmetry, a useful observable-dependent symmetry group may be identified for physically relevant observables, thereby leading to a reduced model of the same size (Sec. VI.1.2). The reduced models may be able to be scaled well-beyond the maximal manageable size of the full model: for example, in some cases we are able to simulate up to 46 bath spins on a typical laptop. We further consider the effect of Markovian dissipation on the bath spins (Sec. VI.1.3), by contrasting the degree of MR attainable under permutation-invariant (collective) noise operators, in which case a strong symmetry is present, with the one attainable if the dissipation is local, in which case only a weak symmetry remains. The reduced models we obtain are refined in Sec. VI.1.4, where we show how further reachable-based MR may be obtained by considering specific initial states.
@@
As a second example, in Sec. VI.2 we analyze a boundary-driven XXZ spin chain. Here, we again demonstrate how observable-based MR can be applied to studying physical quantities of interest to quantum transport, beyond the steady-state regime that is usually considered in the literature. We also again contrast the level of MR that is achievable depending on whether a weak or strong symmetry is present in the dynamics. Finally, in Sec. VI.3 we consider the smallest noiseless-subsystem code for a non-Abelian class of (collective) errors, based on three physical qubits [62], and illustrate how MR can be implemented on the logical level, to obtain a more efficient description of the error dynamics that an encoded qubit undergoes as a result of unwanted error terms. In doing so, we provide another example of MR afforded by an observable-dependent symmetry which is not a (weak) symmetry of the dynamics.
@@
We conclude in Sec. VII by iterating our key findings and their significance, and by discussing some important open questions – along with an outlook to future steps. Technical proofs of mathematical results stated in the main text are provided in Appendix A, whereas Appendixes B and C expand, respectively, on MR derivations and symmetry properties of some of the examples we presented in Sec. VI.
II Quantum semigroups with outputs
II.1 Dynamical settings and problem statement

### v3

@@ Abstract
1 Introduction
1.1 Context and motivation
@@
Motivated by the above applications, a host of strategies for reducing the demands of direct simulation of quantum dynamics and obtaining tractable “effective” models have been developed. As in the classical case, a main guiding principle is the fact that, even if access to a complete description could be granted, this would neither be desirable nor necessary, as only certain information is ultimately “relevant” on physical grounds. Discarding “irrelevant” information can be achieved by invoking different forms of coarse-graining, aggregating, or averaging over microscopic degrees of freedom, typically in combination with perturbative arguments – paradigmatic example being renormalization-group approaches, mean-field and hydrodynamic descriptions, or cluster-expansion techniques [17, 18, 19]. For open quantum systems specifically, weak-coupling and timescale-separation assumptions are used in standard Born-Markov derivations of master equations [20, 2, 21], along with more sophisticated approaches based, for instance, on different forms of clustering or coarse graining in space [22, 23] or time [24, 25]. Further to that, acknowledging that the dynamics of realistic many-body quantum systems (subject to specified locality, energy, and time constraints) may a priori be confined to an exponentially small fraction of the available state space [26] in turn legitimates efficient parametrizations of “relevant” quantum states in terms of tensor networks [27, 19]. The use of tensor-network methods, in combination with chain-mapping techniques [28, 29] or process-tensor representations [30, 31], lies at the core of current state-of-the-art numerical methods available for open quantum system simulation in the presence of highly structured, possibly strongly-coupled environments. As the conceptual and practical significance of open quantum dynamics continues to expand across quantum science, efforts for obtaining improved simulation methods are multiplying, with a special focus on the simplest yet paradigmatic case of dynamics described by time-independent Lindblad master equations [20]. Recent representative contributions include more accurate and efficient classical simulation algorithms based on high-order adiabatic elimination [32, 33] or low-rank structure [34], as well as quantum simulation algorithms that exploit unitary embeddings [35] or classical randomness [36].
@@
From a system-theoretical standpoint, the different approaches mentioned above can all be seen as instances of approximate model (-order) reduction (MR), a general framework aimed at reducing the complexity of a mathematical model while retaining some of the essential features of the original one [37]. A familiar setting in which MR takes place, in fact in its strongest exact form, arises when the dynamics exhibits a manifest symmetry. If the dynamics are described by a Hamiltonian, a decomposition of the Hilbert space according to irreducible representations of the symmetry group naturally furnishes invariant subspaces [38]; hence, if the initial state resides entirely in one such subspace, an exact lower-dimensional Hamiltonian model may be found by restricting to this invariant subspace via projection. This holds true both for “conventional” symmetries, for which the number of dynamically disconnected invariant, “Krylov subspaces” [39, 40] grows at most polynomially with system size, and for non-manifest “unconventional” symmetries, as occurring in systems exhibiting so-called Hilbert space fragmentation [41, 42, 43], whereby exponentially many disconnected Krylov subspaces exist. In this context, an important step was taken by Kumar and Sarovar [44], where conditions for certifying the existence of a non-trivial proper invariant subspace were identified and, building on that, exact MR methods were developed for quantum many-body Hamiltonian models. A subspace of the Hilbert space, however, is not the natural candidate for reduction when the initial condition is not a pure state or the dynamics is not unitary. As such, the approach does not lend itself to the application to genuine open quantum dynamics.
@@
Our framework may describe a many-body open quantum system $S$ evolving on a (finite-dimensional) Hilbert space $\mathcal{H}$ and subject to Markovian dissipation, or it can be applied to reduce the joint (Hamiltonian or Lindblad) dynamics of a many-body bipartite system evolving on $\mathcal{H}\equiv\mathcal{H}_{S}\otimes\mathcal{H}_{E}$, with $E$ being a quantum environment – in which case the reduced dynamics of $S$ alone is generically non-Markovian. While being suitable for describing arbitrary quantum states (including non-factorized system-environment initial conditions), our approach explicitly leverages the fact that, in typical applications, the dynamical behavior of the system is not only needed solely for certain initial conditions, but also for only certain observable properties of interest (e.g., expectations of few-body or collective observables, diagonal elements of the density operator, and so on). This results in the possibility of carrying out dimensional MR based on both initial conditions and output observables, which has not been exploited before to the best of our knowledge. As standard in both quantum probability [47] and rigorous formulations of quantum statistical mechanics [48, 49], we define the dynamical models of interest on operator (*-)algebras. By insisting that projections are also taken onto operator *-subalgebras, we can guarantee that the reduced models we obtain obey the structural constraints of physically admissible (completely positive and trace-preserving, CPTP) quantum dynamics – which we prove are still in Lindblad form. An iterative algorithm is provided to construct a reduced, lower-dimensional Lindbladian model, when one exists. We further show that, even in situations where symmetries are present in the original Lindblad dynamics, our general framework may allow for more effective reduction than achievable by known symmetry-based approaches, through the identification of a new type of observable-dependent symmetries.
@@
Related work. From a technical standpoint, we borrow some key tools and results from our previous work on MR for classical hidden Markov models [50] and quantum discrete-time dynamics [51]. While our guiding philosophy here is similar, extension to continuous-time Markovian dynamics nonetheless entails, as we will see, several non-trivial extensions. In particular, it is by no means obvious – and it is in fact remarkable – that our MR procedures yield a valid Lindblad generator. Further to that, extension to continuous-time dynamics substantially broaden the class of physical settings and phenomena to which our methods may be usefully applied to, given the foundational importance of Lindblad master equations.
@@
Concerning early contributions to the use of MR techniques for quantum dynamics, projection-based methods were employed for nonlinear conditional (stochastic) quantum master equations in the context of cavity QED [52], and later used to exactly derive a special (Maxwell-Bloch-type) class of Lindblad dynamics [53]. In [54], MR by balanced truncation was exploited to derive reduced (exact or approximate) models for linear quantum stochastic dynamics, with emphasis on providing conditions under which physical realizability constraints are preserved. Closer to our approach, both in spirit and in terms of the algebraic formalism employed, are the work by Kumar and Sarovar we mentioned above [44] and the work in [55], where a notion of state-space coarse-graining is formulated for quantum systems. As we already noted, while the approach in [44] also achieves exact quantum MR, a quantum model is obtained “for free” because only Hamitonian and pure-state dynamics are considered. On the flip side, while the approach of [55] is not a priori restricted to Hamiltonian dynamics, the coarse-grained reduced models are not always exact nor provably CPTP or in Lindblad form. More recently, building on the concept of “lumpability” of Markov processes [56], exact quantum MR techniques based on constrained bisimulation have been developed for boosting the simulation of quantum circuits [57]. Aside from the fact that this approach targets discrete-time dynamics (rather than continuous-time as we consider here), as the one in [44] it is again restricted to unitary evolution and pure states in current form.
@@
1.2 Structure and summary of main results
@@
The dynamical models we focus on are formally introduced in Sec. 2, in terms of a Lindblad master equation (in the Schrödinger or dual, Heisenberg picture) paired with a linear output equation – which we term a quantum-dynamical semigroup with output. After exemplifying some basic scenarios of interest that our formalism encompasses, we pose two dimensional MR problems, distinguished by the type of reduced model one is seeking: a reduced CPTP quantum dynamics or, respectively, a general linear dynamical system – in such a way that the specified outputs are exactly reproduced for all times. The essential idea we use for obtaining the desired reduction is also anticipated therein (Sec. 2.2): namely, the reduction of the dynamics to an appropriate Krylov operator subspace.
@@
The next two sections of the paper, Sec. 3 and Sec. 4, are devoted to introduce the mathematical apparatus needed for tackling and solving the MR problems we posed. In particular, the requisite Krylov operator subspaces are built by leveraging the available knowledge of the initial conditions of interest for the system – formalized in the notion of a reachable subspace, introduced in Sec. 3.1 – and the target output observables for the dynamics – formalized in the dual notion of an observable subspace, constructed in Sec. 4.1). In order to extend such subspaces to operator subalgebras, amenable to support quantum dynamics, some existing structural results on operator *-algebras and their associated quantum conditional expectations and “distortion maps” are needed; these are recalled in Sec. 3.2.
In essence, our key results in these theory sections may be summarized as follows.
$\bullet$ Linear reductions. The solution to the linear MR problem is derived in Theorem 1 based on the reachable subspace. Informally, we show that a linear dynamical model of provably minimal dimension and able to reproduce the full state evolution is obtained by “discarding degrees of freedom” that are not reached by the dynamics – formally, by projecting the dynamics onto the Krylov operator subspace generated by the initial conditions. Likewise, by assuming arbitrary initial conditions, a reduced linear dynamical model that is also provably optimal and able to reproduce the full evolution of observables is obtained in Theorem 5 by discarding degrees of freedom that have no observable output – formally, by projecting the dynamics onto the Krylov subspace orthogonal to unobservable operators. These two MRs can be combined together to yield a minimal realization of the dynamics. If $n$ is the dimension of the original state space, the reduced model can be obtained with a single run of an algorithm whose complexity scales may be shown to scale roughly as $n^{8}$.
$\bullet$ Quantum reductions. The solution to the quantum MR problem requires, as mentioned, that the reduced dynamics be themselves defined on an operator *-algebra. We obtain such algebras by essentially “closing”, in an appropriate mathematical sense, the reachable and observable subspaces we mentioned above. If the relevant *-algebra, say, $\mathscr{A}$, is smaller than the full operator algebra associated to the system, the key feature we leverage to achieve MR stems from a well-known structural characterization that $\mathscr{A}$ enjoys, the Wedderburn decomposition [62, 63]. Pictorially, the idea is illustrated in Fig.  1: since every *-algebra can be seen as the direct sum of possibly repeated irreducible blocks, a smaller-dimensional representation may be obtained by “compressing” $\mathscr{A}$ into a reduced *-algebra, say, $\check{\mathscr{A}}$, for which no repeated block appears. In Theorem 4, we formally prove that a Lindblad generator ${\cal L}$
@@
retains its CPTP properties when appropriately restricted to an operator subalgebra in this way, and a reduced Lindblad generator $\check{\cal L}$ may be obtained from ${\cal L}$ through composition with suitable “reduction” and “injection” maps. A constructive algorithm is proposed, to compute the desired reduction starting from the specified set of initial conditions. Notably, the initial states can belong to an arbitrary linear set, without the need to assume initial factorization in the case a composite system-environment setting, as in standard derivations of master equations [2]. In Sec. 4.2, we derive an equivalent result for reductions based on the observable set of interest, and the two procedures are combined in an iterative algorithm described in Sec. 4.3. In a worst-case scenario, each run of this algorithm still scales as $O(n^{8})$.
Figure 1: Pictorial illustration of the basic principle enabling quantum MR. The Wedderburn decomposition of a representative operator *-algebra, $\mathscr{A}=(\mathbb{1}_{2}\otimes\mathbb{C}^{4\times 4})\bigoplus(\mathbb{1}_{3}\otimes\mathbb{C}^{2\times 2})\bigoplus\mathbb{C}^{3\times 3}\subseteq\mathbb{C}^{17\times 17}$, is depicted in the top, while its compressed representation, $\check{\mathscr{A}}=\mathbb{C}^{4\times 4}\bigoplus\mathbb{C}^{2\times 2}\bigoplus\mathbb{C}^{3\times 3}\subseteq\mathbb{C}^{9\times 9}$, is given at the bottom. The action of the reduction and injection maps, $\mathcal{R}$ and $\mathcal{J}$, is also shown, with formal definitions given in Sec. 3.2 and 3.3.
@@
Section 5 bridges between the general theory sections and the illustrative applications that will follow, by focusing on the interplay of our general MR procedures with symmetries which, for Lindblad dynamics, are usually categorized as “strong” or “weak” – depending on whether, loosely speaking, invariance occurs at the level of individual (Hamiltonian and Lindblad) operators or only at the superoperator level [64, 65]. We show how viable subspaces and algebras for MR naturally emerge from considering symmetries, and characterize the sense in which a strong symmetry leads to a computational advantage over a weak one, from a dimensional-reduction standpoint. Nonetheless, focusing on observable-based MR, we prove in Theorem 7 that our methods can improve reductions based on both strong and weak symmetries. In fact, finding the relevant output algebra by the procedure described above corresponds to finding “generalized symmetries” that emerge only when considering the observables of interest: this new class of observable-dependent symmetries is defined in Sec. 5.4.
@@
In Sec. 6, we examine concrete examples motivated by condensed-matter physics and quantum information science applications, with a twofold goal: on the one hand, to benchmark our general approach, by showing that it consistently recovers known exact results when available, or it yields the correct numerical results of the full model, as expected; on the other hand, to showcase additional flexibility and resource reduction, beyond what achievable with existing methods to the best of our knowledge. Specifically, in Sec. 6.1 we study variations of the well-known central-spin model, beginning with the case where the dynamics is Hamiltonian (no dissipation) and enjoys permutation symmetry on the bath spins. Here, we observe how the introduction of a bath Hamiltonian $H_{B}$ influences the achievable MR, by showing how, even when $H_{B}$ breaks permutation symmetry, a useful observable-dependent symmetry group may be identified for physically relevant observables, thereby leading to a reduced model of the same size (Sec. 6.1.2). The reduced models may be able to be scaled well-beyond the maximal manageable size of the full model: for example, in some cases we are able to simulate up to 46 bath spins on a typical laptop. We further consider the effect of Markovian dissipation on the bath spins (Sec. 6.1.3), by contrasting the degree of MR attainable under permutation-invariant (collective) noise operators, in which case a strong symmetry is present, with the one attainable if the dissipation is local, in which case only a weak symmetry remains. The reduced models we obtain are refined in Sec. 6.1.4, where we show how further reachable-based MR may be obtained by considering specific initial states.
@@
As a second example, in Sec. 6.2 we analyze a boundary-driven XXZ spin chain. Here, we again demonstrate how observable-based MR can be applied to studying physical quantities of interest to quantum transport, beyond the steady-state regime that is usually considered in the literature. We also again contrast the level of MR that is achievable depending on whether a weak or strong symmetry is present in the dynamics. Finally, in Sec. 6.3 we consider the smallest noiseless-subsystem code for a non-Abelian class of (collective) errors, based on three physical qubits [66], and illustrate how MR can be implemented on the logical level, to obtain a more efficient description of the error dynamics that an encoded qubit undergoes as a result of unwanted error terms. In doing so, we provide another example of MR afforded by an observable-dependent symmetry which is not a (weak) symmetry of the dynamics.
@@
We conclude in Sec. 7 by iterating our key findings and their significance, and by discussing some important open questions – along with an outlook to future steps. Technical proofs of mathematical results stated in the main text are provided in Appendix A.1, whereas considerations on the implementation complexity of our MR procedures are given in Appendix A.2. Appendixes B and C expand, respectively, on MR derivations and symmetry properties of some of the examples we presented in Sec. 6.
2 Quantum semigroups with outputs
2.1 Dynamical settings and problem statement